OpenStudy (anonymous):

Point C is graphed on a number line at –7. Point D is 13 units away from point C on the number line. What could be the coordinates of point D? Choose all answers that are correct. A. –20 B. 3 C. 6 D. 20

OpenStudy (ashes.boom):

basically, count 13 spaces from negative 7

OpenStudy (ashes.boom):

-7 + 13 and -7 - 13 those are the two answers you will have

OpenStudy (ashes.boom):

what do you think? @vmeza0131

OpenStudy (anonymous):

first one is 6 the second one is -20

OpenStudy (anonymous):

@Ashes.boom

OpenStudy (ashes.boom):

yup :) nice job

OpenStudy (anonymous):

can you help me with some more @Ashes.boom

OpenStudy (ashes.boom):

yea sure whats up

OpenStudy (anonymous):

What value of z makes the statement true? –(z + (–17)) = 12 + 17 A. –46 B. –12 C. 12 D. 46

OpenStudy (anonymous):

@Ashes.boom

OpenStudy (ashes.boom):

okay so the way i would do this is simplify it as -z - 17 = 12 + 17 then you want to get the variable on one side of the equation by itself. What would you do next?

OpenStudy (ashes.boom):

@vmeza0131

OpenStudy (anonymous):

@Ashes.boom i got -12

OpenStudy (ashes.boom):

okay how did you do it

OpenStudy (ashes.boom):

I got something different

OpenStudy (anonymous):

−(z−17)=12+17−(z−17)=12+17 −z+17=12+17(Distribute) −z+17=(12+17)(Combine Like Terms) −z+17=29 −z+17=29Step 2: Subtract 17 from both sides. −z+17−17=29−17 −z=12Step 3: Divide both sides by -1. −z −1 = 12 −1 z=−12 Answer: z=−12

OpenStudy (anonymous):

@Ashes.boom

OpenStudy (ashes.boom):

okay yea that works lol idk what i did wrong but i just got -12 :)

OpenStudy (anonymous):

can you help me with one more @Ashes.boom

OpenStudy (ashes.boom):

yea shoot :)

OpenStudy (anonymous):

On a number line, point S has coordinate –1.6 and point T has coordinate –6.5. Which expressions represent the distance between point S and point T? Choose all answers that are correct.

OpenStudy (ashes.boom):

what are the answer choices??

OpenStudy (anonymous):

1.|1.6+(-6.5)| 2.|-1.6-(-6.5)| 3.|-6.5-(-1.6)| 4.|-6.5-1.6|

OpenStudy (ashes.boom):

okay whacha thinking

OpenStudy (anonymous):

so i have to choose all that are correct so i think its 1 and 4 but i dont know @Ashes.boom

OpenStudy (ashes.boom):

okay one sec ill be right back

OpenStudy (ashes.boom):

Okay i just checked 1 and 4 and they do not come out as the same answer

OpenStudy (anonymous):

i don't remember how to do this @Ashes.boom

OpenStudy (ashes.boom):

I might be totally wrong but i think 1, 2, and 3 are all right

OpenStudy (ashes.boom):

the way i did it was drew out a number line and graphed the two points. Then i counted from -1.6 to -6.5. I got 4.9 spaces. Since all of the equations above have the absolute value symbols it means the negatives and positives are neutral. the second one comes out as -4.9 but its absolute value is 4.9 which matches 1 and 3 so i say all three (1, 2, and 3)

OpenStudy (ashes.boom):

make sense??

OpenStudy (ashes.boom):

i gotta go but i hope that helped

OpenStudy (anonymous):

that did bye